OpenStudy (anonymous):

The length of a rectangle is 7 inches more than 2 times the width. If the area of the rectangle is 114 square inches, find the length and the width. ?????????????

OpenStudy (lgbasallote):

let w = width length is 7 inches more than twice the width so l = 2w + 7 area is 114 so (2w + 7)w = 114 => 2w^2 + 7w = 114 => 2w^2 + 7w - 114 = 0 solve for w

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

-19/2, 6 ???

OpenStudy (lgbasallote):

width can't be negative so you just consider 6

OpenStudy (anonymous):

yes i got it thank you so much!!

OpenStudy (lgbasallote):

welcome